OpenStudy (anonymous):

|x - 7| - 7 = -3 how do i answer this?

OpenStudy (jadeishere):

Simplify |x - 7| - 7 = - 3 Simplify |* Simplify -7|-7= Simplify -3 -3 is a constant, so we do not do anything with this Our expression becomes | + - 3| Our answer is -3

OpenStudy (jadeishere):

Sorry confusing, but -7 = 7 so -7 + 7 = 0 while x stays x so x = -3

OpenStudy (anonymous):

Thanx @Jadeishere

OpenStudy (jadeishere):

yep :)

OpenStudy (the_fizicx99):

Wut?

OpenStudy (jadeishere):

Lol xD

OpenStudy (the_fizicx99):

Dafuq, Add 7 to both sides, its absolute value so its the distance from that point to 0. You have 2 equations, one will be negative, the other positive. |x - 7| - 7 = -3 +7 +7 |x - 7| = 4 and |x - 7| = -4 Then solve by dropping the absolute values signs and adding 7 to both sides

OpenStudy (jadeishere):

absolute value isn't my best subject so you are most likely right

OpenStudy (the_fizicx99):

o.o